# GUIDE: Guidance-based Incremental Learning with Diffusion Models

### 06 Mar 2024 v1

@@ D.1 Coverage of data manifold
In this analysis, we explore the forgetting behavior of a diffusion model trained continually on CIFAR-10, divided into two tasks with 25000 samples each. Initially, we train the model exclusively on data from the first task (Real Task 1). Subsequently, we employ two approaches: standard self-rehearsal training for the second task (Continual Task 2), akin to our method, and retraining the first task’s model on the entire CIFAR-10 dataset (Upper-bound Task 2), which serves as our upper bound. All training conditions, including architectures, training steps (100k), and hyperparameters, remain consistent across setups. Generative metrics (FID (Heusel et al., 2017), Precision, and Recall (Kynkäänniemi et al., 2019)) for samples generated with DDIM100 from all models are displayed in Tab. 5.
Continual training of the diffusion model leads to a noticeable reduction in its capability to cover the training manifold of Task 1, as evidenced by a decrease in the Recall metric. However, the Precision metric does not show a significant loss, indicating that the quality of the generated samples remains largely unaffected. Moreover, we visualize the umap embeddings of both real data examples and generated samples in Fig. 8. In plot (c), it’s evident that the coverage of real data samples (depicted in blue) by the generated samples has noticeably diminished.
D.2 Forgetting of training data samples
In our study, we further analyze what data samples the diffusion model trained continually tends to forget. We conducted an experiment where we sample examples from class apple from our diffusion model, trained on the CIFAR-100 dataset divided into 5 tasks, after completing the second task. Utilizing Precision and Recall metrics definitions (Kynkäänniemi et al., 2019), we computed representations using the Inception-v3 model for both real and generated sample sets. This allowed us to identify which segments of the approximated real data manifold were covered or missed by the generated manifold.
Figure 7(b) illustrates examples of both remembered and forgotten sets. Consistent with findings by Toneva et al. (2018), our diffusion model tends to forget rare samples in the training set, such as those with complex backgrounds, while retaining more common samples characterized by simpler backgrounds and typical shapes. This is also connected to the drop in Recall, where from task to task, the diffusion model loses the ability to generate samples from low-density regions of data manifold (Sehwag et al., 2022).
@@ D.2 Forgetting of training data samples
Table 5: Comparison of FID, Recall, and Precision metrics of all three trained models. When we train the diffusion model in a continual manner we observe a very significant drop in Recall.
@@
(c) Initial manifold Task 1.
(d) Manifold of Task 1 after training on Task 2 - upper-bound.
(e) Manifold of Task 1 after training on Task 2 - continual.
@@
Figure 8: Visualization of data manifold of Task 1 in CIFAR-10 dataset. Blue points represent embedded real data samples from Task 1, and orange points represent generated samples. We see a significant drop in coverage of training data manifold after training on the second task of continual training.
@@
Figure 9: Random samples of airplanes (left) and automobiles (right) from our unconditional diffusion model trained only on airplanes and automobiles from the CIFAR-10 dataset. Samples are generated using 1000 denoising steps. We use guidance from two classifier models - one guides each sample to either automobile or airplane class, while the other one guides to the class denoted in each row. Specifically, in the first row, we do not use the guidance from the second classifier.

### 31 May 2024 v2

@@ E.1 Coverage of data manifold
In this analysis, we explore the forgetting behavior of a diffusion model trained continually on CIFAR-10, divided into two tasks with 25000 training samples each. Initially, we train the model exclusively on data from the first task (Real Task 1). Subsequently, we employ two approaches: standard self-rehearsal training for the second task (Continual Task 2), akin to our method, and retraining the first task’s model on the entire CIFAR-10 dataset (Upper-bound Task 2), which serves as our upper bound. All training conditions, including architectures, training steps (100K), and hyperparameters, remain consistent across setups. Generative metrics (FID [16], Precision, and Recall [27]) for samples generated with DDIM100 from all models are displayed in Tab. 9.
@@
Table 9: Comparison of FID, Recall, and Precision metrics of all three trained models. We calculate metrics for 25000 samples generated using DDIM100. When we train the diffusion model in a continual manner, we observe a very significant drop in the Recall.
@@
Continual training of the diffusion model leads to a noticeable reduction in its capability to cover the training manifold of Task 1, as evidenced by a decrease in the Recall metric. However, the Precision metric does not show a significant drop, indicating that the quality of the generated samples remains largely unaffected. Moreover, we visualize the umap embeddings of both real data examples and generated samples in Fig. 8. In plot (c), it’s evident that the coverage of real data samples (depicted in blue) by the generated samples has noticeably diminished.
(a) Initial task 1 data manifold.
(b) Task 1 data manifold after training on task 2 – upper-bound.
(c) Task 1 data manifold after training on task 2 – continual training
@@
Figure 8: Visualization of task 1 data manifold in the CIFAR-10 dataset. Blue points represent embedded real data samples from task 1, and orange points represent generated samples. We see a significant drop in coverage of training data manifold after training on the second task of continual training.
E.2 Forgetting of training data samples
In our study, we further analyze what data samples the diffusion model trained continually tends to forget. We conducted an experiment where we sample examples from class apple from our class-conditional diffusion model, trained on the CIFAR-100 dataset divided into 5 tasks, after completing the second task. Utilizing Precision and Recall metrics definitions [27], we computed representations using the Inception-v3 model for both real and generated sample sets. This allowed us to identify which segments of the approximated real data manifold were covered or missed by the generated manifold.
Figure 9(a) illustrates examples of remembered data samples, while Figure 9(b) depicts those data samples that were forgotten by the diffusion model due to continual training. Consistent with findings by Toneva et al. [44], our diffusion model tends to forget rare samples in the training set, such as those with complex backgrounds, while retaining more common samples characterized by simpler backgrounds and typical shapes. This is also connected to the drop in Recall, where from task to task, the diffusion model loses the ability to generate samples from low-density regions of data manifold [40].
Appendix F Runtime analysis of GUIDE for CIFAR-10/5
We perform an extensive runtime analysis in order to compare the training times of GUIDE to other existing baselines. For the training of each method, we used the same machine with a single NVIDIA RTX A5000 GPU. Since in the training of GUIDE on CIFAR-10/5 setup, we generate new rehearsal examples every 10 batches, for fairness, we evaluate the DDGR runtime with exactly the same interval. We present the calculated runtimes in Tab. 10.
Table 10: Runtime analysis of all baseline methods on the CIFAR-10/5 benchmark. (*) We evaluate the DDGR method with the same rehearsal generation interval as we use in GUIDE on this benchmark – once every 10 batches.
Method Time [GPU-hours]
DGR VAE 1.28
DGR+distill 1.35
RTF 1.09
BIR 2.49
GFR 1.76
DDGR $\text{111.96}^{*}$
DGR diffusion 76.1
GUIDE 82.95
Approaches that train diffusion models (DDGR, DGR diffusion, and GUIDE) require significantly more computations. However, GUIDE does not increase much of the runtime compared to the standard DGR with diffusion model ($+9\%$ wall time). At the same time, when comparing GUIDE to the recently proposed DDGR [12] method that also uses diffusion models in generative replay scenario, our approach needs notably less computation time ($-26\%$ wall time) while achieving better performance in all evaluated setups.
Appendix G Speed vs accuracy trade-off in GUIDE
We investigate the trade-off between speed and performance of our method. To that end, we study the effect of the number of denoising steps used for rehearsal sampling and the replay’s set size on the performance of GUIDE. In both experiments, we present average accuracy after the final task, along with a speed-up in time.
To investigate the effect of the number of denoising steps used for rehearsal sampling, we test the performance of GUIDE for $[10,20,50,100,250,1000]$ steps of DDIM (Tab. 11). Although the performance tends to increase with a number of denoising steps, the difference in average accuracy is not very significant.
To effectively reduce the size of the replay set, we generate new rehearsal samples once every $[1,5,10,50,100]$ batches of the training (Tab. 12). For completeness, we also present the results for fine-tuning setup - training without any generative replay that can be seen as an upper bound in terms of computational complexity. We observe that increasing the generation interval can not only significantly decrease the training time but also increase the final performance.
Although the sampling process of diffusion models is indeed relatively slow compared to other generative architectures, we show that we can significantly reduce the training time of GUIDE without a notable drop in performance by using either one of the proposed techniques.
Table 11: Trade-off between the number of denoising steps used for the rehearsal sampling on the effectiveness of GUIDE on CIFAR-10/5 benchmark (default setup from the main text is grayed out).
DDIM steps $\bar{A}_{T}$ Time [speedup]
10 $59.7\pm 1.04$$\times$ 7.3
20 $63.7\pm 0.99$$\times$ 2.3
50 $64.1\pm 0.09$$\times$ 1
100 $64.2\pm 0.54$$\times$ 0.51
250 $63.6\pm 0.25$$\times$ 0.21
1000 (no DDIM) $71.8\pm 0.23$$\times$ 0.06
Table 12: Trade-off between the interval of the generation of new rehearsal samples on the effectiveness of GUIDE on CIFAR-10/5 benchmark (default setup from the main text is grayed out).
Rehearsal generation interval $\bar{A}_{T}$ Time [speed-up]
1 $63.12\pm 0.45$$\times$ 0.2
5 $64.47\pm 0.45$$\times$ 1
10 $64.1\pm 0.09$$\times$ 1.95
50 $59.2\pm 0.47$$\times$ 8.49
100 $56.9\pm 0.98$$\times$ 14.41
$\infty$ (Fine-tuning) $18.95\pm 0.20$$\times$ 49.59
Appendix H Samples from ImageNet100 128x128
Figure 10: Random samples generated from an unconditional diffusion model trained only on goldfish and tiger shark classes from the ImageNet100 128x128 dataset. Each image grid presents samples generated with guidance to the class depicted above the grid which is unknown to the diffusion model. We generate samples using 1000 denoising steps and we set guidance scale $s$ to $10$ for both classes that we guide to.
Appendix I Samples from CIFAR-10
@@
Figure 11: Samples from unconditional diffusion model trained only on classes automobile and airplane from the CIFAR-10 dataset. We generate those samples following the same procedure as discussed in Sec. 4.1. On the left grid, we guide the denoising process to automobiles, while on the right grid, to airplanes. At the same time, we add the guidance to classes depicted on the y-axis of the figure. We generate samples using 1000 denoising steps and we set guidance scale $s$ to $5$ for both classes that we guide to.
Appendix J Broader impacts
This paper presents work whose goal is to advance the field of continual machine learning. There are many potential societal consequences of our work, many of which are generic to the machine learning field in general, i.e., the DGR algorithm considered in this paper will reflect the biases present in the dataset. Hence, it is important to exercise caution when using this technique in applications where dataset biases could lead to unfair outcomes for minority and/or under-represented groups. In our case, this especially concerns training a diffusion model used for replay, in which simple random sampling can result in a different data distribution than the original dataset. It would be worthwhile to actively monitor the model’s outputs for fairness or implement bias correction techniques to mitigate these negative impacts.